# Amortized Bayesian Workflow

### v2

@@ C.4 Decision model
Following von Krause et al. [89], we specify the prior distributions for the drift-diffusion model parameters as:
@@
$a_{1},a_{2}$ $\sim\text{Gamma}(6,0.15),$
$\tau_{c}$ $\sim\text{Gamma}(3,0.15),$
$\tau_{n}$ $\sim\text{Gamma}(3,0.5),$
where all Gamma distributions use the shape–scale parametrization. [^8] The prior distributions for the boundary separation parameters $a_{1}$ and $a_{2}$ differ slightly from those in von Krause et al. [89] due to a different parametrization of boundary separation. We implement the drift-diffusion model likelihood using the hssm package [21] and PyMC.
@@
We use $100,000$ simulated parameter–observation pairs for training the amortized approximator, $1000$ for validation, and $200$ for training-phase diagnostics, including parameter recovery and simulation-based calibration.
@@
The closed-world diagnostics (parameter recovery and simulation-based calibration checking) in Figure 10 indicate that the neural network training has successfully converged to an acceptable posterior approximator within the scope of the training set.
@@
The test datasets consist of 15,000 participants pre-processed from the online implicit association test (IAT) database [96, 89]. Each test dataset is a tensor of shape $(120,4)$, where each row corresponds to a single trial and contains the response time, missing data mask, experiment condition type, and stimulus type.
Appendix D Amortized initialization for NUTS
In addition to ChEES-HMC, we evaluate the effectiveness of amortized posterior draws as initializations for the NUTS sampler. As shown in Figure 11, amortized initializations reduce the number of required warm-up iterations for both the GEV problem and the decision model. For the psychometric curve and Bernoulli GLM problems, all three initialization methods (amortized, PSIS-refined, and random) yield similar convergence behavior according to $\hat{R}$ diagnostic.
@@ Appendix D Amortized initialization for NUTS
Figure 11: The effect of initialization for NUTS. The figure shows median±$\pm$±IQR across 20 test datasets. Using amortized posterior draws as initializations for ChEES-HMC reduces the required warmup in the GEV and decision model tasks.

### v3

@@ C.4 Decision model
Following von Krause et al. (2022), we specify the prior distributions for the drift-diffusion model parameters as:
@@
[…]
where all Gamma distributions use the shape–scale parametrization. [^10] The prior distributions for the boundary separation parameters $a_{1}$ and $a_{2}$ differ slightly from those in von Krause et al. (2022) due to a different parametrization of boundary separation. We implement the drift-diffusion model likelihood using the hssm package (Fengler et al., 2025) and PyMC.
@@
We use $100,000$ simulated parameter–observation pairs for training the amortized estimator, $1000$ for validation, and $200$ for training-phase diagnostics, including parameter recovery and simulation-based calibration.
@@
The closed-world diagnostics (parameter recovery and simulation-based calibration checking) in Figure 10 indicate that the neural network training has successfully converged to an acceptable posterior estimator within the scope of the training set.
@@
The test datasets consist of 15,000 participants pre-processed from the online implicit association test (IAT) database (Xu et al., 2014; von Krause et al., 2022). Each test dataset is a tensor of shape $(120,4)$, where each row corresponds to a single trial and contains the response time, missing data mask, experiment condition type, and stimulus type.
Appendix D Additional experimental study of the OOD diagnostic in Step 1
To further investigate the relationship between the Mahalanobis distance in the OOD diagnostic and the quality of the amortized posterior, we visualize this relationship using scatter plots in 11(a) for the four tasks considered in the main text. For each task, we use around 1000 test datasets and compute the Pearson correlation coefficient $r$. The Mahalanobis distance is positively correlated with the two posterior quality metrics (W1 and MMTV) for the GEV, psychometric curve, and decision model tasks, where out-of-distribution test datasets are present. For the Bernoulli GLM, the correlation is negative; here, all test datasets were generated from the same distribution (prior simulations) as the training datasets and the Mahalanobis distance is not informative. From 11(a), we see a key limitation of the Step-1 OOD diagnostic: the Mahalanobis distance is clearly not a perfect proxy for the posterior quality. In particular, the amortized estimator may still yield low-quality posterior draws on a dataset with a smaller Mahalanobis distance, as also observed in Figure 5.
We next check the impact of the threshold $\alpha$ in the Step-1 OOD diagnostic by varying it from 0.01 to 0.5, specifically over the set $[0.01,0.05,0.1,0.15,0.2,0.3,0.4,0.5]$, as shown in 11(b). As $\alpha$ increases, more test datasets are rejected, and the overall posterior quality of accepted amortized posterior draws generally improves as measured by the median and IQR of the posterior metrics (lower posterior metric values indicate higher quality). The quality of rejected amortized posterior draws also improves as $\alpha$ increases, while remaining consistently worse than that of the accepted amortized draws. [^11] For Bernoulli GLM, where test datasets and training datasets come from the same distribution, we observe a slightly reverse trend as $\alpha$ increases. This is consistent with the corresponding result shown in 11(a) and further suggests that the Mahalanobis distance is not a good measure for posterior quality for in-distribution datasets in this case. Overall, the posterior quality of the accepted amortized draws, in terms of median and IQR of W1 and MMTV, is not very sensitive to the threshold $\alpha$, and $\alpha=0.05$ appears to be a reasonable default choice.
These results support the use of the Mahalanobis-distance-based OOD test as a lightweight first-line diagnostic in Step 1: it tends to flag the most problematic datasets and thereby improves the quality of accepted amortized posterior draws at negligible additional cost. At the same time, the residual low-quality posteriors at small Mahalanobis distances underscore that this diagnostic cannot guarantee accuracy. For applications that require tighter accuracy guarantees, it is therefore natural to enforce escalation to Step 2 (PSIS) irrespective of the OOD outcome, trading additional computation for a more robust posterior approximation.
(a) Posterior quality metrics (W1 and MMTV) versus Mahalanobis distance for the four benchmark tasks.
(b) Sensitivity of the rejection threshold $\alpha$ in the OOD test. The median $\pm$ IQR (shaded area) of the posterior quality metrics is shown separately for accepted and rejected datasets.
Figure 11: Relationship between amortized posterior quality metrics, Mahalanobis distance, and the OOD rejection threshold in Step 1. (a) Scatter plots of W1 and MMTV versus Mahalanobis distance, with Pearson correlation coefficient $r$ reported in each panel. (b) Effect of varying the threshold $\alpha$ on the posterior quality of accepted and rejected amortized posterior draws. See text in Appendix D for details.
Appendix E Amortized initialization for NUTS
In addition to ChEES-HMC, we evaluate the effectiveness of amortized posterior draws as initializations for the NUTS sampler. The experimental settings mirror those used for ChEES-HMC (Section 3.4), except that we launch only four chains, which is the typical configuration for NUTS. As shown in Figure 12, amortized initializations reduce the number of required warm-up iterations for both the GEV problem and the decision model. For the psychometric curve and Bernoulli GLM problems, all three initialization methods (amortized, PSIS-refined, and random) yield similar convergence behavior according to the $\hat{R}$ diagnostic (Vehtari et al., 2021).
@@ Appendix E Amortized initialization for NUTS
[…]